Question 1 [4 marks]

A proton is accelerated through a potential difference of 3 kV, and then enters a magnetic field of flux density 0.6 T, in a direction perpendicular to the direction of the magnetic field. It travels at right angles to the magnetic field at all times. If the charge-to-mass ratio for the proton is 9.53x107 C kg-1, calculate the radius of the proton’s circular orbit.
1
Expert's answer
2021-09-13T11:26:36-0400

"V= 3 \;kV = 3000 \;V \\

B = 0.6 \;T \\

\frac{g}{m} = 9.53 \times 10^7 \;c/kg"

Radius of proton’s circular orbit

"r=\frac{1}{B} \times \sqrt{\frac{2mV}{q}} \\

= \frac{1}{B} \times \sqrt{\frac{2V}{q/m}} \\

r = \frac{1}{0.6} \times \sqrt{\frac{2 \times 3000}{9.53 \times 10^7}} \\

= 1.3224 \times 10^{-2} \;m \\

= 1.32 \;cm"